Examine the intelligence quotient, or IQ, as it is defined by the formula:
“intelligence quotient = 100 * (mental age/chronological age)”
Justify why it is reasonable for the mean to be 100.
Please can someone walk me through this

Answer by stanbon(75887)

Examine the intelligence quotient, or IQ, as it is defined by the formula:
“intelligence quotient = 100 * (mental age/chronological age)”
Justify why it is reasonable for the mean to be 100.
-----
If your mental age is equal to your chronological age you are
an average person, neither smarter or dumber than other people.
----
In that case your IQ = 100*(your mental age your chron age) = 100*1 = 100
-------------------
Cheers,
Stan H.